OpenStudy (anonymous):

Nitrogen has two naturally occurring isotopes, N-14 and N-15. the atomic mass of nitrogen is 14.007 amu. which isotope is more abundant in nature? Explain

OpenStudy (anonymous):

sleek-feathered one

OpenStudy (anonymous):

wut

OpenStudy (anonymous):

first open textbook and read what is an isotope, then you will understand that isotopes difffer by neutron number and if as you say there are N-14 and N-15 isotopes of N and its atomic mass is 14.007 then it becomes clear that more aboundant is N-14 for if it were lets say 50% N-14 and 50% of N-15 its atomic mass would be 14.5, so from 14,007 you can calculate that there is only 0,05% of N-15 ( x(N15) = (0,0007/14,007) * 100% = 0,05%)